# A Comprehensive Survey on Kolmogorov Arnold Networks (KAN)

### v6

@@ 6.1 Time Series
TKAN (Temporal Kolmogorov-Arnold Networks)
@@
T-KAN and MT-KAN (Time Series Kolmogorov-Arnold Networks and Multivariate Time Series Kolmogorov-Arnold Networks)
@@
SigKAN (Signature-Weighted Kolmogorov-Arnold Networks)
@@ 6.2 Interpretability and Symbolic Regression
MonoKAN (Monotonic Kolmogorov-Arnold Network)
@@
BKANs (Bayesian Kolmogorov Arnold Networks )
@@ 6.6 Physics-Informed Neural Networks and Engineering Applications
KINN (Kolmogorov-Arnold-Informed Neural Network)
@@
PIKAN (Physics-Informed Kolmogorov-Arnold Networks)
@@ 7.6 Theoretical Analysis
Gao and Tan [65] studied gradient descent (GD) and stochastic gradient descent (SGD) of KAN. They claim that using SGD to optimize KAN can be used to enhance the performance on some machine learning tasks. Moreover, they analysis the global convergence of GD and SGD for physics-informed KANs, which might be used to optimize physics-informed KANs. Wang et al [66] carefully compared KAN and MLP from a theoretical perspective. From the perspective of approximation, MLP and KAN is very similar sharing parameterize similar classes of functions. However, they find that KANs lack the spectral bias toward low frequencies that is characteristic of MLPs. They claim that their research results can contribute to computing applications like solving PDEs.

### v7

@@ 6.1 Time Series
Temporal Kolmogorov-Arnold Networks (TKAN)
@@
Time Series Kolmogorov-Arnold Networks and Multivariate Time Series Kolmogorov-Arnold Networks (T-KAN and MT-KAN)
@@
Signature-Weighted Kolmogorov-Arnold Networks (SigKAN)
@@ 6.2 Interpretability and Symbolic Regression
Monotonic Kolmogorov-Arnold Network (MonoKAN)
@@
Bayesian Kolmogorov Arnold Networks (BKANs)
@@ 6.4 Computer Vision and Image Processing
PEP-GS
Jin et al [47] combine Multi-head Self-attention, KAN and Laplacian Pyramid Decomposition and propose a new framework PEP-GS. They claim that their framework show high accuracy in complex scenarios.
@@ 6.6 Physics-Informed Neural Networks and Engineering Applications
Kolmogorov-Arnold-Informed Neural Network (KINN)
@@
Physics-Informed Kolmogorov-Arnold Networks (PIKAN)
@@
ChebPIKAN
Based on Physics-informed Neural Networks (PINNs) and Chebyshev polynomial, Guo et al [54] propose a new framework ChebPIKAN. They show the strong ability of the model in solving PDEs through numerous experiments. This model demonstrates strong potential in fields requiring complex and rapid computations like Fluid Mechanics.
Separable Physics-Informed Kolmogorov-Arnold Networks (SPIKANs)
In order to solve the problem of slow training speed, Jacob, Howard and Stinis [55] introduce a new framework which combines Separable PINNs and KAN. The special design of the framework can separate the PDEs problems into different parts and use KAN payers to solve rather than traditional unstructured point-cloud used in PINNs. These improvements lead to a reduction in the time required to solve PDE problems.
6.7 Tabular Data
TKGMLP
Zhang et al [56] propose a hybrid model called TKGMLP, which combines MLP and KAN and it is designed for large-scale financial tabular data. Moreover, they propose a new encoding method for financial numerical data. They claim that their model can enhance the prediction accuracy for tasks such as credit scoring.
KACDP (Kolmogorov - Arnold Credit Dedault Predict)
Liu and Zhao [57] use KAN for personal credit risk prediction. They claim that their model has a better performance compared with traditional models. They thought that the non-linear property of KAN is the key and the model has great interpretability.
@@ 7.6 Theoretical Analysis
Gao and Tan [70] studied gradient descent (GD) and stochastic gradient descent (SGD) of KAN. They claim that using SGD to optimize KAN can be used to enhance the performance on some machine learning tasks. Moreover, they analysis the global convergence of GD and SGD for physics-informed KANs, which might be used to optimize physics-informed KANs.
Wang et al [71] carefully compared KAN and MLP from a theoretical perspective. From the perspective of approximation, MLP and KAN is very similar sharing parameterize similar classes of functions. However, they find that KANs lack the spectral bias toward low frequencies that is characteristic of MLPs. They claim that their research results can contribute to computing applications like solving PDEs.
@@ 7.7 Improvements and Optimizations
Kolmogorov-Arnold Networks with Interactive Convolutional Elements (KANICE)
Ferdaus et al [80] combine the Interactive Convolutional Blocks (ICBs) of RNN and KAN linear layers with CNN framework, which can be used to computer vision field. Compared with traditional frameworks, they show that their model is more accurate. Also, they propose a new version of KANICE with less parameters but has almost the same accuracy.